If m1,m2 and m3 be the masses of three particles v12, v23 and v13 be their relative velocities, prove that the total kinetic energy (E) of the system about the centre of mass is given by

m1m2v212+ m2m3v223+ m1m3v213 / m1+ m2 + m3

1
Expert's answer
2021-09-04T15:19:31-0400

"\vec v_{12}=\vec{v_1}-\vec{v_2}"

"\vec v_{23}=\vec{v_2}-\vec{v_3}"

"\vec v_{13}=\vec{v_1}-\vec{v_3}"


"m_1m_2\vec v_{12}^2+m_2m_3\vec v_{23}^2+m_1m_3\vec v_{13}^2="

"m_1m_2(\vec{v_1}-\vec{v_2})^2+m_2m_3(\vec{v_2}-\vec{v_3})^2+m_1m_3(\vec{v_1}-\vec{v_3})^2="

"\text{(After expanding the brackets and grouping the terms )}"

"(m_1\vec v_1^2+m_2\vec v_2^2+m_3\vec v_3^2)(m_1+m_2+m_3)-"

"(m_1\vec v_1+m_2\vec v_2+m_3\vec v_3)^2\ (1)"

"\text{Divide the resulting expression (1) by }(m_1+m_2+m_3)"

"\text{We also take into account the fact that:}"

"T= \frac{m_1\vec v_1^2}{2}+ \frac{m_2\vec v_2^2}{2}+ \frac{m_3\vec v_3^2}{2}"

"(m_1\vec v_1^2+m_2\vec v_2^2+m_3\vec v_3^2)= 2T"

"\text{where }T - \text{total kinetic energy of the system}"


"(m_1\vec v_1+m_2\vec v_2+m_3\vec v_3)^2"

"P = m_1\vec v_1+m_2\vec v_2+m_3\vec v_3"

"\text{where } P\text{ center of gravity impulse}"

"T_c =\frac{1}{2}(m_1+m_2+m_3)v_c^2"

"T_c -\text{kinetic energy of the center of gravity}"

"P =(m_1+m_2+m_3)v_c"

"(m_1\vec v_1+m_2\vec v_2+m_3\vec v_3)^2= 2(m_1+m_2+m_3)T_c"


"\frac{m_1m_2\vec v_{12}^2+m_2m_3\vec v_{23}^2+m_1m_3\vec v_{13}^2}{m_1+m_2+m3}=2T-2T_c"

"\text {by Koenig's theorem}"

"T = T_c+T_r"

"T_r -\text{ relative kinetic energy of the system}"


"\frac{m_1m_2\vec v_{12}^2+m_2m_3\vec v_{23}^2+m_1m_3\vec v_{13}^2}{m_1+m_2+m3}=2T_r"

"\text{hence the hypothesis is false}"


"\text{Answer: the proposed proof formula is false}"